1. Calculate the internal energy and enthalpy of 1 kg of air occuping 0.05 m (cubic) at 20 bar. If the internal energy is increased by 120 kJ as the air is compressed to 50 bar, calculate the new volume, Cp and Cv.
1
Expert's answer
2020-09-16T14:42:28-0400


Assuming air to be a calorically perfect gas.

R=287J/Kg-K, Cp=1.005KJ/kgK

Cv=0.714 KJ/kgK

Given as: mass(m)=1Kg, Volume (V)=0.05m2

By ideal gas equation:

"\frac{P \times V }{Rm}=T"

T=(20x105x0.05)/287 =348.43 K

Internal energy (U1)=mCvT=1x714x3484.3=248.780 k J

If internal energy is increased by 120 kJ

U2=248.78+120=368.78kJ

T2=U2/Cv=516.5K

Given as P2=50x105Pa

"\frac{P \times V }{Rm}=T"

V2=(mRT2)/(P2)

=(287x0.05x516.5)/(50x105)

=0.001467 m3